OpenStudy (anonymous):

Medals will be given need help with problems, have the answers already just need to work them out. 1.x^2 + 20x + 99 2.n^2-13n+ 40 3. x^2-8x-20 5. x^2 + 7xy - 44y^2 6. 2x^2 + 7x +6

OpenStudy (anonymous):

are u looking for the factors?

OpenStudy (anonymous):

no just to work the problem out i already have the answer

OpenStudy (anonymous):

1.1*99=99 11*9=99 11+9=20 2.1*40=40 -8*-5=40 -8+(-5)=-13 3.1*-20=-20 -10+2=-8 -10*2=-20 4.1*-44=-44 11-4=7 11*-4=-44 5.2*6=12 4+3=7 4*3=12

OpenStudy (anonymous):

What's the name of this lesson? Like, is it Radical expressions or sumthin like dat?

OpenStudy (anonymous):

@surjithayer can you please number them and space them ? its confusing

OpenStudy (anonymous):

only three rows for every problem. i have given hint only to make factors.

OpenStudy (anonymous):

are you sure this is right ? i have the answers

OpenStudy (anonymous):

yes ,absolutely correct.

OpenStudy (anonymous):

Do your answers match his?

OpenStudy (anonymous):

do you have #6 ?

OpenStudy (anonymous):

now i figured 4 th is missing. so consider 4th as 5th and 5 as 6th.

OpenStudy (anonymous):

@Goku-Kai he didnt give me the answers he just showed the work i dont need the answers!!!

OpenStudy (anonymous):

@bettylee No, I thought you said you have the answers already and he worked through the problem so I didn't know if y'all answers matched up that's all. I'm sorry...

OpenStudy (anonymous):

x^2+7xy-44y^2 =x^2+(11-4)xy-44y^2 =x^2+11xy-4xy-44y^2 =x(x+11y)-4y(x+11y) =(x+11y)(x-4y)

OpenStudy (anonymous):

no i didnt say it rudely lol i was just saying i dont want anyone giving me the answers because i have them i justb really need to work the problem out loll

OpenStudy (anonymous):

@surjithayer thank you so much really helpful i have some more if you want to help if not then ill just post them

OpenStudy (anonymous):

Oh lol, It's hard to comprehend the emotions of people by reading text lol. But I gotchu, it's all love.

OpenStudy (anonymous):

lollll its okay it can be sometimes i looked back at the message and it does sound rude but it wasnt lmaooo. all love!!!

OpenStudy (anonymous):

when the product of co-efficient of x^2 and constant term is positive,then break the coefficient of x in two parts such that their product is same . when product is negative then coefficient of x=larger -some number such their product is same and difference is coefficient of x.

OpenStudy (anonymous):

okay i understand thank you for explaining that for me really you were a big help!!! @surjithayer

OpenStudy (anonymous):

yw